# RoboDuet: Whole-body Legged Loco-Manipulation with Cross-Embodiment Deployment

### v4

@@ VI Discussion and Limitations
In this letter, we propose RoboDuet, a whole-body mobile manipulation framework that integrates two collaborative policies for velocity tracking and 6D end-effector pose control, enabling agile and robust performance across diverse tasks. Leveraging a two-stage training approach, RoboDuet effectively utilizes robust locomotion priors to enhance the system’s resistance to external disturbances. Additionally, our method supports cross-embodiment deployment, allowing seamless hardware replacement without retraining, and it achieves real-world performance comparable to simulation.
While RoboDuet focuses on low-level whole-body control, achieving fully autonomous task execution will require integration with high-level planners and advanced collision detection. These are key areas for future enhancement.

### v5

@@ VI Discussion and Limitations
In this letter, we propose RoboDuet, a whole-body legged loco-manipulation framework that integrates two collaborative policies for velocity tracking and 6D end-effector pose control, enabling agile and robust performance across diverse tasks. Leveraging a two-stage training approach, RoboDuet effectively utilizes robust locomotion priors to enhance the system’s resistance to external disturbances. Additionally, our method supports zero-shot transfer deployment, allowing seamless hardware replacement without retraining, and it achieves real-world performance comparable to simulation.
While RoboDuet focuses on low-level whole-body control, achieving fully autonomous task execution will require integration with high-level planners and advanced collision detection. Moreover, our method can only handle terrains such as slopes and gravel while maintaining a fixed target pose. However, it has yet to achieve loco-manipulation on more complex terrains, such as stairs. These are key areas for future enhancement.